OpenStudy (exoexoexo):

a weight of 120 pounds is placed five feet from the fulcrum of a lever. how far from the fulcrum should a 100 pound weight be placed in order to balance the lever?

Vocaloid (vocaloid):

(weight 1)(distance 1) = (weight 2)(distance 2) (120)*5 = (100)*(distance 2) distance 2 =?

OpenStudy (exoexoexo):

6 feet

Vocaloid (vocaloid):

great, good job!

OpenStudy (exoexoexo):

thank you